OpenStudy (anonymous):

Help!! will give medal:) :) how do I find the scale factor for the dilation of ΔABC into ΔDEF

OpenStudy (anonymous):



OpenStudy (ankit042):

can you find lenght of CB and FE use the distance formula

OpenStudy (anonymous):

For vertical lines only the difference of y co-ordinates \[\frac{ DE }{AB }=\frac{ 6-(-3) }{4-(-2) }=\frac{ 9 }{6 }=\frac{ 3 }{2 }\]

OpenStudy (anonymous):

so the scale factor is 3/2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright thanks so much :)